OpenStudy (vflies):

A scale drawing of a kitchen is shown below. The scale is 1 : 20. A rectangle is shown. The length of the rectangle is labeled 4 inches. The width of the rectangle is labeled 5 inches. Show your work to determine the area of the room in square feet.

OpenStudy (wolf1728):

Will you be posting a graphic?

OpenStudy (wolf1728):

Okay, we don't need a graphic.

OpenStudy (kinged):

Multiply both dimensions by 20 and do the rectangle area formula

OpenStudy (kinged):

Then convert your answer to feet

OpenStudy (vflies):

4*20=80 and 5*20 = 100

OpenStudy (vflies):

@Kinged

OpenStudy (wolf1728):

Image is 4 inches by 5 inches = .3333333 by 0.416666 feet We have to increase these by 20 so .3333333 * 20 by 0.416666 * 20 feet = 6.6666666667 feet by 8.3333333333 feet I think you can calculate the area now.

OpenStudy (kinged):

What's even easier wold be converting the dimensions to feet before doing the formula.

OpenStudy (wolf1728):

Kinged which I did about 2 seconds ago

OpenStudy (vflies):

I don't understand

OpenStudy (kinged):

Which part?

OpenStudy (vflies):

So I multiplied the sides by 20 Then I convert

OpenStudy (kinged):

Yep. That's also right.

OpenStudy (vflies):

So wolf converted it for me what do I do next?

OpenStudy (vflies):

@Kinged @wolf1728

OpenStudy (vflies):

@Kinged

OpenStudy (kinged):

Do the rectangle formula.

OpenStudy (vflies):

What are they?

OpenStudy (vflies):

What do we have to do to calculate the formula?

OpenStudy (wolf1728):

area = length times width

OpenStudy (vflies):

What do we have to do to calculate the formula?

OpenStudy (wolf1728):

What is the width and what is the length?

OpenStudy (vflies):

So 4*5=20

OpenStudy (wolf1728):

That is the area of the image

OpenStudy (vflies):

Ok

OpenStudy (vflies):

I don't know how to calculate the formula

OpenStudy (wolf1728):

I gave you the dimensions in feet 6.6666666667 feet by 8.3333333333 feet

OpenStudy (vflies):

So 55.555556

OpenStudy (wolf1728):

Yes that is it!!

OpenStudy (wolf1728):

Now here's an answer I've been working on for a while. I think the "tricky" part of this question is the conversion from square inches to square feet For example, you could say the image is 5 by 4 inches. Image area is 20 square inches. (correct) Now here's the "tricky part: You then increase it by 20 and assume that the actual size is 400 square inches, which it is incorrect. Since we are dealing with area, we have to increase our answer by a factor of 20 TIMES 20. So, our answer is 20 square inches * 20 TIMES 20 = 8,000 square inches There are 12 inches per foot and 12 TIMES 12 square inches per square foot or 144 square inches per square foot So, 8000 square inches divided by 144 = 55.5555555556 square feet

OpenStudy (vflies):

Are we done?

OpenStudy (wolf1728):

Yes we are done. I posted that long explanation for other folks that might look at this problem.

OpenStudy (kinged):

That's the best question ever, baby.

OpenStudy (kinged):

You can go play, now.

OpenStudy (vflies):

Thanks Guys

OpenStudy (wolf1728):

u r welcome

OpenStudy (vflies):

Thanks Guys